OpenStudy (anonymous):

solve by the substition , 9m+n=15 m-9n=29

OpenStudy (danmac0710):

m=29+9n Sub into first equation: 9(29+9n)+n=15 261+81n+n=15 261+82n=15 82n=15-261 n=-246/82 n=3 Is that okay?

OpenStudy (anonymous):

so it would be 246/82 would be your solution